OpenStudy (hlilly2413):

Common table sugar, sucrose, has a MM of 342.3 g/mol. To what volume (in mL) should 10.0g of sucrose be diluted to give a solution containing 0.500 moles/liter? I understand that the mass is 10.0g, the molar mass is 342.3g/mol, and that I need to find the volume in mL, but am unsure of how to go about it. Please help.

OpenStudy (jhannybean):

Use the Molarity as a conversion factor to find L of solution. \[\sf 10.0~g ~ C_6H_{12}O_6~ \times ~ \frac{1 ~mol~ C_6H_{12}O_6}{342.3~g~ C_6H_{12}O_6}~ \times \frac{1~L~soln}{0.500~mol~C_6H_{12}O_6}~\times \frac{1~mL}{10^{-3}~L}\]\[\sf =\boxed{58.4 ~mL~ soln}\]

OpenStudy (jhannybean):

sorry, to find mL of solution*